# Stock Movement Prediction with Multimodal Stable Fusion via Gated Cross-Attention Mechanism

### v1

@@ 4.4 Fine-grained movement prediction
$\textbf{h}_{i,d,g}=MLP_{t}(\textbf{H}_{i,d,g}),\textbf{h}_{i}=MLP_{t}(\textbf{H}_{i})$ (17)
@@
5 Performance evaluation
@@
    Prediction performance: Does MSGCA outperform previous methods in multimodal stock movement prediction?
@@
    Ablation study: How do fusion strategies, language encoders, and graph encoders affect the performance of MSGCA?
@@
    Modality effect: How do different modalities affect the performance of MSGCA? Which modality contributes the most?
@@
    Stable fusion study: Does the gated cross-attention mechanism in MSGCA work to stabilize multimodal fusion?
@@
    Hyperparameter effect: How do hidden dimension sizes, time window sizes and learning rates affect MSGCA performance?
@@
    Computation efficiency: How does the training speed and memory usage of MSGCA comparing to baseline methods?
5.1 Experimental setup
1) Datasets: We evaluate the performance of MSGCA on three public benchmark datasets: BigData22 [16], ACL18 [10], CIKM18 [59], and one new dataset that we collect and publish: InnoStock. Table 3 presents a summary of these datasets, detailing the count of stocks, the number of news documents (InnoStock) and tweet documents (others), the number of edges in the industry graph and the duration of trading dates. The three existing datasets consist of high-trade-volume stocks in US stock markets and dynamic input from tweets and price sequences. Innostock, originally collected by us from CSMAR  [^4] https://cn.gtadata.com/, focuses on newly formed technology companies listed on China’s Sci-Tech Innovation Board, aggregating their financial news from various online platforms. To accommodate MSGCA’s multimodal input, we further collect the industrial sector relationships for each stock from the above datasets to build graphs. We label each stock according to the increase rate of its adjusted closed prices. Following [16], the increase rate is calculated as $r_{t}^{s}=p_{t}^{s}/p_{t-1}^{s}-1$, where $p_{t}^{i}$ is the adjusted closing price of the stock $s$ at the timestamp $t$. To ensure balanced labels in our fine-grained classification, we adjust the growth rate threshold for flat trends according to each dataset. We set $[-1\%,1\%]$, $[-0.5\%,0.5\%]$, $[-0.4\%,0.4\%]$, and $[-0.3\%,0.3\%]$ for InnoStock, BigData22, ACL18, and CIKM18, respectively. For example, InnoStock timestamps that have $r_{t}^{s}\geq 1\%$ and $r_{t}^{s}\leq-1\%$ are labeled as up (1) and down (-1), respectively, and $-1\%<r_{t}^{s}<1\%$ are labeled as flat (0). We chronologically partitioned each dataset into training, validation, and testing subsets, consistent with recent studies on stock movement prediction [16] [36].
Table 3: Summary of datasets statistics
@@ 5.2 Baseline methods
5.3 Performance comparison (RQ1)
Stock movement prediction performance are evaluated with the two metrics. The results are shown in Table 4, where the best results are highlighted in bold, and the most competitive baselines are underlined. Our MSGCA consistently achieves the best performance on all datasets and metrics. Specifically, we have the following observations.
Table 4: Results of stock movement prediction on four datasets and two metrics
Methods InnoStock BigData22
ACC MCC ACC MCC
LSTM 0.3545±0.0005 0.0320±0.0011 0.4119±0.0023 0.0866±0.0059
ALSTM 0.3762±0.0017 0.0339±0.0014 0.4335±0.0079 0.0968±0.0072
DTML 0.4208±0.0009 0.0493±0.0013 0.4367±0.0123 0.1048±0.0131
ESTIMATE 0.3995±0.0020 0.4820±0.0046 0.4296±0.0027 0.1004±0.0059
ALSTM-W 0.3835±0.0015 0.0431±0.0022 0.4287±0.0025 0.0999±0.0066
SLOT 0.4127±0.0006 0.0509±0.0005 0.4250±0.0082 0.0980±0.0081
MSGCA 0.4223±0.0003 0.0550±0.0008 0.4379±0.0077 0.1112±0.0037
Methods ACL18 CIKM18
ACC MCC ACC MCC
LSTM 0.3779±0.0022 0.0254±0.0020 0.3781±0.0006 0.0393±0.0028
ALSTM 0.3806±0.0027 0.0275±0.0031 0.3830±0.0011 0.0470±0.0046
DTML 0.3832±0.0053 0.0306±0.0085 0.3811±0.0080 0.0613±0.0072
ESTIMATE 0.3805±0.0046 0.0312±0.0074 0.3806±0.0047 0.0579±0.0035
ALSTM-W 0.3848±0.0050 0.0375±0.0064 0.3818±0.0082 0.0545±0.0096
SLOT 0.3894±0.0016 0.0487±0.0168 0.3848±0.0011 0.0601±0.0053
MSGCA 0.3966±0.0050 0.0593±0.0073 0.3861±0.0023 0.0807±0.0045
@@ 5.4 Ablation study (RQ2)
5.5 Effect of multimodal sources (RQ3)
In this subsection, we examine how various modalities aid in predicting stock movement. To evaluate this, we implement three variants of MSGCA where a modality is removed from our trimodal encoder module: 1) MSGCA-ID: a version in which indicator sequences and dynamic documents are considered without a relational graph as input; 2) MSGCA-IG: a version that indicator sequences and graph are kept for feature fusion; 3) MSGCA-DG: a version using only documents and relationships, excluding the primary indicator modality. The performance of these variants on four datasets is shown in Table 5. We can observe from the results that MSGCA-ID and MSGCA-IG still outperform existing stock movement prediction methods on most metrics, demonstrating the capability of our feature fusion and movement prediction modules. For example, the MCC of MSGCA-ID is 1.7% and 9.8% higher than that of SLOT on InnoStock and BigData22, respectively. In addition, the modality of indicator sequences plays a crucial role in the prediction performance. All results of MSGCA-DG decrease dramatically after the indicator modality is removed from those of other frameworks. Furthermore, without relying on indicator features, MSGCA can yield results compared to baseline methods that use indicators alone, showing its capability to effectively utilize the information from graph relationships and the textual data of news and tweets for prediction.
Table 5: Effects of different modalities on stock movement prediction
@@ 5.5 Effect of multimodal sources (RQ3)
5.6 Study for stable fusion (RQ4)
@@
Figure 6: Stable fusion study by reducing embedding dimensions of a stock to a single dimension, allowing for direct comparison with its price trend. the top blue line represents the price sequence of the stock. Yellow lines are features after cross-attention fusion for aggregating document and graph modalities without guidance by the primary modality. Green lines are features from gated cross-attention blocks, exhibiting smoother fluctuations than yellow lines and closer to the indicator sequence, indicating the stable fusion effect.
5.7 Effect of hyperparameters (RQ5)
@@
5.8 Computation efficiency analysis (RQ6)
@@
6 Conclusion and future work
@@
[…]
Acknowledgements This work is supported by the Key Research and Development Program of Zhejiang Province, China (No. 2021C01013).
[…]
Author contributions Chang Zong: Conceptualization, Software, Validation, Writing. Jian Shao: Methodology, Investigation. Weiming Lu: Methodology, Validation. Yueting Zhuang: Review, Editing.
[…]
Data availability The four benchmark datasets used in our study are available from our Google Drive at https://drive.google.com/drive/folders/1Tqtq2kpkHt6RXSiLWSql1fgOwsSyZdIL?usp=sharing.
[…]
Ethical and informed consent for data used This article does not contain any studies with human participants or animals performed by any of the authors.
[…]
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
ORCID Chang Zong https://orcid.org/$0$-0001-7757-0659

### v2

@@ IV-D Fine-Grained Movement Prediction
$\textbf{h}_{i,d,g}=MLP_{t}(\textbf{H}_{i,d,g})$ (21)
$\textbf{h}_{i}=MLP_{t}(\textbf{H}_{i})$ (22)
@@
IV-E Pseudocode of Training MSGCA
To conclude the training process of MSGCA with the above phases, the pseudocode is illustrated as Algorithm 1.
Algorithm 1 Pseudocode of Training MSGCA
Input: Indicator sequence: =$\mathcal{I}=\{i_{1},i_{2},...,i_{t}\}$, Dynamic documents: $\mathcal{D}=\{d_{1},d_{2},...,d_{t}\}$, Relational graph: $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{U}\}$
Arguments:
Hidden dimension: $d$, Window size: $ws$, Epoch number: $en$, Batch size: $bs$, Learning rate: $lr$
Output: Training loss: $\mathcal{L}$, Model parameters: $\boldsymbol{\theta}$
1:Get embeddings of dynamic documents $\textbf{v}_{d}^{{}^{\prime}}$ from LLM
2:Get indicator sequences $\textbf{s}_{c}$, $\textbf{s}_{o}$, $\textbf{s}_{h}$
3:Split instances according to window size $ws$
4:Sample batched training set $\mathcal{B}$ according to batch size $bs$
5:for $epoch\in 1,...,en$ do
6:     for $batch\in\mathcal{B}$ do
7:         Calculate indicator features $\textbf{v}_{i}$ using Eq. 1-4
8:         Calculate document features $\textbf{v}_{d}$ using Eq. 5-6
9:         Calculate graph features $\textbf{v}_{g}$ using Eq. 7-9
10:         Get fused features $\textbf{H}_{i,d}^{l}$ from indicator and document modalities using Eq. 10-11
11:         Get stable features $\textbf{H}_{i,d}$ using Eq. 12-14
12:         Get fused features $\textbf{H}_{i,d,g}^{l}$ from previous features and graph modality using Eq. 15-16
13:         Get stable features $\textbf{H}_{i,d,g}$ using Eq. 17-19
14:         Get aggregated features h using Eq. 20-22
15:         Get probability vectors p using Eq. 23
16:         Calculate multi-label loss $\mathcal{L}$ using Eq. 24
17:         Return $\mathcal{L}$
18:     end for
19:     Update $\theta$ with backpropagation from loss $\mathcal{L}$ using Adam optimizer with learning rate $lr$
20:end for
21:Return model parameter $\boldsymbol{\theta}$
V Experiments
@@
    Prediction performance (Section V-C): Does MSGCA outperform previous methods in multimodal stock movement prediction?
@@
    Ablation study (Section V-D): How do fusion strategies, language encoders, and graph encoders affect the performance of MSGCA?
@@
    Modality effect (Section V-E): How do different modalities affect the performance of MSGCA? Which modality contributes the most?
@@
    Stable fusion study (Section V-F): Does the gated cross-attention mechanism in MSGCA work to stabilize multimodal fusion?
@@
    Hyperparameter effect (Section V-G): How do hidden dimension sizes, time window sizes and learning rates affect MSGCA performance?
@@
    Computation efficiency (Section V-H): How does the training speed and memory usage of MSGCA comparing to baseline methods?
V-A Experimental Setup
1) Datasets: We evaluate the performance of MSGCA on three public benchmark datasets: BigData22 [13], ACL18 [7], CIKM18 [57], and one new dataset that we collect and publish: InnoStock. Table III presents a summary of these datasets, detailing the count of stocks, the number of news documents (InnoStock) and tweet documents (others), the number of edges in the industry graph and the duration of trading dates. The three existing datasets consist of high-trade-volume stocks in US stock markets and dynamic input from tweets and price sequences. Innostock, originally collected by us from CSMAR  [^5] https://cn.gtadata.com/, focuses on newly formed technology companies listed on China’s Sci-Tech Innovation Board, aggregating their financial news from various online platforms. To accommodate MSGCA’s multimodal input, we further collect the industrial sector relationships for each stock from the above datasets to build graphs. We label each stock according to the increase rate of its adjusted closed prices. Following [13], the increase rate is calculated as $r_{t}^{s}=p_{t}^{s}/p_{t-1}^{s}-1$, where $p_{t}^{i}$ is the adjusted closing price of the stock $s$ at the timestamp $t$. To ensure balanced labels in our fine-grained classification, we adjust the growth rate threshold for flat trends according to each dataset. We set $[-1\%,1\%]$, $[-0.5\%,0.5\%]$, $[-0.4\%,0.4\%]$, and $[-0.3\%,0.3\%]$ for InnoStock, BigData22, ACL18, and CIKM18, respectively. For example, InnoStock timestamps that have $r_{t}^{s}\geq 1\%$ and $r_{t}^{s}\leq-1\%$ are labeled as up (1) and down (-1), respectively, and $-1\%<r_{t}^{s}<1\%$ are labeled as flat (0). We chronologically partitioned each dataset into training, validation, and testing subsets, consistent with recent studies on stock movement prediction [13] [33].
TABLE III: Summary of Datasets Statistics.
@@ V-B Baseline Methods
V-C Performance Comparison (RQ1)
Stock movement prediction performance are evaluated with the two metrics. The results are shown in Table IV, where the best results are highlighted in bold, and the most competitive baselines are underlined. Our MSGCA consistently achieves the best performance on all datasets and metrics. Specifically, we have the following observations.
TABLE IV: Results of Stock Movement Prediction on Four Datasets and Two Metrics.
Methods InnoStock BigData22 ACL18 CIKM18
ACC MCC ACC MCC ACC MCC ACC MCC
LSTM 0.3545±0.0005 0.0320±0.0011 0.4119±0.0023 0.0866±0.0059 0.3779±0.0022 0.0254±0.0020 0.3781±0.0006 0.0393±0.0028
ALSTM 0.3762±0.0017 0.0339±0.0014 0.4335±0.0079 0.0968±0.0072 0.3806±0.0027 0.0275±0.0031 0.3830±0.0011 0.0470±0.0046
DTML 0.4208±0.0009 0.0493±0.0013 0.4367±0.0123 0.1048±0.0131 0.3832±0.0053 0.0306±0.0085 0.3811±0.0080 0.0613±0.0072
ESTIMATE 0.3995±0.0020 0.4820±0.0046 0.4296±0.0027 0.1004±0.0059 0.3805±0.0046 0.0312±0.0074 0.3806±0.0047 0.0579±0.0035
ALSTM-W 0.3835±0.0015 0.0431±0.0022 0.4287±0.0025 0.0999±0.0066 0.3848±0.0050 0.0375±0.0064 0.3818±0.0082 0.0545±0.0096
SLOT 0.4127±0.0006 0.0509±0.0005 0.4250±0.0082 0.0980±0.0081 0.3894±0.0016 0.0487±0.0168 0.3848±0.0011 0.0601±0.0053
MSGCA 0.4223±0.0003 0.0550±0.0008 0.4379±0.0077 0.1112±0.0037 0.3966±0.0050 0.0593±0.0073 0.3861±0.0023 0.0807±0.0045
@@ V-D Ablation Study (RQ2)
V-E Effect of Multimodal Sources (RQ3)
In this subsection, we examine how various modalities aid in predicting stock movement. To evaluate this, we implement three variants of MSGCA where a modality is removed from our trimodal encoder module: 1) MSGCA-ID: a version in which indicator sequences and dynamic documents are considered without a relational graph as input; 2) MSGCA-IG: a version that indicator sequences and graph are kept for feature fusion; 3) MSGCA-DG: a version using only documents and relationships, excluding the primary indicator modality. The performance of these variants on four datasets is shown in Table V. We can observe from the results that MSGCA-ID and MSGCA-IG still outperform existing stock movement prediction methods on most metrics, demonstrating the capability of our feature fusion and movement prediction modules. For example, the MCC of MSGCA-ID is 1.7% and 9.8% higher than that of SLOT on InnoStock and BigData22, respectively. In addition, the modality of indicator sequences plays a crucial role in the prediction performance. All results of MSGCA-DG decrease dramatically after the indicator modality is removed from those of other frameworks. Furthermore, without relying on indicator features, MSGCA can yield results compared to baseline methods that use indicators alone, showing its capability to effectively utilize the information from graph relationships and the textual data of news and tweets for prediction.
TABLE V: Effects of Different Modalities on Stock Movement Prediction.
@@ V-E Effect of Multimodal Sources (RQ3)
V-F Study for Stable Fusion (RQ4)
@@
Figure 6: Stable fusion study by reducing embedding dimensions of a stock to a single dimension, allowing for direct comparison with its price trend. Blue lines represent the price time-series of the stock. Yellow lines are features after cross-attention fusion for aggregating document and graph modalities without guidance by the primary modality. Green lines are features from gated cross-attention blocks, exhibiting smoother transitions than yellow lines and indicating stable fusion effects.
V-G Effect of Hyperparameters (RQ5)
@@
V-H Computation Efficiency Analysis (RQ6)
@@
VI Conclusion and Future Work
@@
Acknowledgment
We would like to thank the anonymous reviewer for his/her careful reading of our manuscript and his/her many insightful comments and suggestions.